1.determine the area and show this graphically





A. To the right of z=1.3





B. Between z=-1 and z=-2.3






2.find each of the following percentile points under the normal curve complete your procedure





A. P79





B. P90

1
Expert's answer
2022-03-23T04:59:33-0400

1) A."P(Z>1.3)=1-P(Z<1.3)=0.90320=0.0968"

B."P(-2.3<Z<-1)=P(Z<-1)-P(Z<-2.3)=0.15866-0.01072=0.14794"




2) A. "P(Z<a)=0.79\implies a=0.807"

B. "P(Z<a)=0.9\implies a=1.282"